the sum of two angle is 130. if the larger is 10 more than twice the smaller find the 2 angle

Answer by simranb(63)

Let the smaller angle be x
then larger angle=2x+10
their sum=130
so
x+2x+10=130
3x+10=130
3x=130-10
3x=120
x=120/3
x=40
then 2x+10=2*40+10
=80+10=90
so the two angles are 90 and 40
Hope this helps!